Flux F = \frac{energy}{(unit area) ∗ (unit time)}

= \frac{E}{dA ∗ dt}

F_{λ}dλ = \frac{E_{λ}dλ}{dA ∗ dt ∗ dλ}

I_{λ} = \frac{E_{λ}dλ}{dA ∗ dt ∗ dλ ∗ cos θ ∗ dΩ}

u_{λ}dλ = \frac{1}{c}∫ I_{λ}dλdΩ

u_{λ}dλ = \frac{4π}{c} I_{λ}dλ

u_{λ}dλ = \frac{8πhc/λ^{5}}{e^{(hc/λkT)} − 1}dλ

u_{λ}dλ = \frac{4σT^{4}}{c}

p = \frac{E}{c}

P_{rad}dλ = \frac{1}{c}∫ I_{λ}dλ cos^{2} θdΩ

P_{rad}dλ = \frac{4π}{3c} I_{λ}dλ

P_{rad}dλ = \frac{4π}{3c} ∫ B_{λ}(T)dλ = \frac{4σT^{4}}{3c} = \frac{1}{3}u